### 23 Mar 2026 v2

# Confidential Databases Without Cryptographic Mappings

Abstract
Confidential databases (CDBs) are essential for enabling secure queries over sensitive data in untrusted cloud environments using confidential computing hardware. While adoption is growing, widespread deployment is hindered by high performance overhead from frequent synchronous cryptographic operations, which causes significant computational and memory bottlenecks.
We present Fedb, a novel CDB design that removes cryptographic operations from the critical path. Fedb leverages crypto-free mappings, which maintain data-independent identifiers within the database while securely mapping them to plaintext secrets in a trusted domain. This paradigm shift reduces the runtime overhead by up to 78.0 on industry-standard benchmarks including TPC-C and TPC-H.
1 Introduction
Cloud databases have become the backbone of modern digital infrastructure, with organizations increasingly outsourcing their data management to Database-as-a-Service (DBaaS) providers [1, 2, 3]. However, this convenience comes with a critical security risk: cloud administrators have unrestricted access to all tenant data, leaving opportunities for both malicious and accidental exposure. Recent breaches, such as the 2025 Allianz Life exposure of 1.1 million personal records [4], underscore this threat. For sensitive databases regulated by laws such as HIPAA [5] and PCI-DSS [6], entrusting cloud infrastructure with plaintext data is untenable.
Confidential databases (CDBs) have emerged as a compelling solution. By leveraging trusted execution environments (TEEs), CDBs enable secure data upload and query execution while ensuring data remains encrypted from the cloud provider. Over the past two decades, researchers have explored various CDB designs [7, 8, 9, 10, 11], and major cloud providers have deployed CDB systems [12, 13, 14, 15, 16].
Despite this industry momentum, a critical roadblock to widespread adoption of CDB systems remains: severe performance degradation compared to plaintext databases, with analytical queries experiencing up to 79.5 higher latency [15]. This gap arises from a deliberate system design choice made by industry deployments: rather than enclosing the entire DBMS within a TEE, which would prevent database administrators from performing essential maintenance tasks, modern CDBs execute only critical operators (e.g., numeric SUM operations) inside TEEs. The resulting split architecture necessitates synchronous software-based decryption and encryption for every remote procedure call (RPC) across domains, as illustrated in Figure 1. The computational burden scales linearly with data volume. For instance, a SUM aggregation over an -row table requires 2-2 decryptions and -1 encryptions in total. Even worse, in order to support arbitrary computation across fields, modern CDBs employ field-level encryption, inflating each 4-byte integer to 32 bytes (8 expansion) due to cryptographic metadata (12-byte nonce and 16-byte authentication tag required by schemes like AES-GCM). This dramatically amplifies cross-domain data movement, memory pressure, and I/O bandwidth demands. These bottlenecks raise a fundamental question: Is cryptographic overhead inherent to CDB design, or can we achieve the same security guarantees with far cheaper alternatives?
Key insight. We argue that the performance crisis stems from an unexamined assumption that has persisted for decades: any data leaving TEEs must be encrypted. Modern CDBs store ciphertexts in the DBMS and decrypt them in TEEs for computation. We observe that this is fundamentally a mapping-based protection scheme: the DBMS uses ciphertexts as pointers to reference plaintext data in TEEs. Current practice conflates indirection with protection, encrypting both pointers and data. We decouple these concerns: pointers use lightweight indirection; only data at rest uses encryption. Based on this insight, we aim to replace the costly ciphertext-to-plaintext mappings with highly efficient index-to-plaintext mappings, while preserving data confidentiality and database correctness.
| Modern CDBs | Fedb | |
| Indirection | Ciphertexts | Data-independent FIDs |
| Protection | Ciphertexts | Encryption at rest |
| RPC cost | Decrypt (1,500 cycles) + Encrypt (5,000 cycles) | Get/Put (50 cycles) |
Our approach. We present Fedb, a CDB system that eliminates compute-intensive cryptography from the critical path without sacrificing security. Fedb introduces crypto-free mappings that bind data-independent field identifiers (FIDs) to their corresponding plaintext fields within TEEs. This allows the DBMS to manipulate data as before, but replaces slow per-field de/encryption with ultra-low-latency put/get operations, as compared in Table 1. In addition, to avoid encrypting every field synchronously, Fedb defers encryption: when data must be evicted from TEE memory to untrusted storage, Fedb performs asynchronous block-level encryption rather than fine-grained field-level encryption. In short, crypto-free mappings deliver two key benefits: they remove cryptographic operations from query execution hot paths and reduce metadata overhead via coarser-grained encryption.
We evaluate Fedb against HEDB [15], a state-of-the-art modern CDB that achieves the trifecta of functionality, maintainability, and security. Our results show that Fedb achieves substantial performance improvements while preserving all desired properties. On the transactional benchmark TPC-C, Fedb achieves up to 1.8 speedup over HEDB. On the analytical benchmark TPC-H, Fedb reduces query execution time by up to 78.0. Fedb outperforms HEDB with 2.9-4.9 speedups. Additionally, Fedb reduces storage overhead by 80.0% and 34.3% across two workloads, respectively. Fedb has been open-sourced and integrated into production GaussDB systems111https://gitee.com/openeuler/secGear/pulls/354.
Contributions. We make the following contributions:
We identify and analyze the performance bottlenecks of modern CDBs through comprehensive profiling and empirical analysis.
We propose a novel CDB design with crypto-free mappings, which remove costly cryptographic operations from the critical path of existing modern CDBs.
We build and deploy Fedb, outperforming state-of-the-art CDBs across transactional and analytical benchmarks.
2 Background and Motivation
2.1 Confidential Computing
Confidential computing [17, 18, 19, 20, 21, 22, 23, 24, 25, 26] provides hardware-based trusted execution environments (TEEs) with three security guarantees. First, secure isolation through either enclaves [18, 23, 22] or confidential virtual machines [17, 19, 20] protects against inspection or tampering by untrusted privileged software. Second, hardware root-of-trust enables clients to verify code authenticity via remote attestation. Third, memory encryption and optional integrity checks resist physical attacks. Confidential computing is now widely available on commercial clouds [27, 28, 21, 29, 30].
2.2 Confidential Databases
Database-as-a-Service (DBaaS) [31] offers fully managed databases that handle updates, backups, and maintenance for users. However, since cloud environments are inherently untrusted, sensitive data like healthcare and financial information requires secure storage and processing.
Confidential databases (CDBs) [7, 8, 9, 32, 10, 11, 15, 12, 14, 13, 16] address this by leveraging confidential computing to enable SQL query processing over encrypted data without compromising confidentiality. Industry-deployed CDBs [12, 14, 13, 15] adopt a split architecture: only essential expression operators (e.g., arithmetic, comparison222All operators yield encrypted results, except for comparisons (e.g. , , ), which output plaintext boolean values for use in indexing ciphertexts.) execute inside TEEs, while most DBMS logic (e.g., SQL parsing, planning, buffer management) remains in the untrusted domain. This minimizes the TCB and preserves maintainability: DBAs can inspect query plans and debug DBMS issues. We refer to these industry-adopted systems as “modern CDBs.”
To enable flexible computation across arbitrary fields, modern CDBs apply field-level encryption. For example, since all fields from the column Salary are encrypted, "SELECT SUM(Salary) FROM Employee" requires the DBMS to iteratively call the addition operator inside the TEE, which decrypts operands, computes the sum, re-encrypts the result, and returns it to DBMS. While secure, this design incurs substantial performance overhead, as we will analyze later.
HEDB [15]: a representative modern CDB. Modern CDBs could be vulnerable to smuggle attacks [15]: malicious DBAs exploit operator interfaces to construct known ciphers (via arithmetic operators, e.g., division yields encrypted “one,” addition builds sequences) and compare them against victim ciphertexts (via equality operators), thereby extracting secrets without authorization. HEDB [15] addresses smuggle attacks through a dual-TEE architecture: an integrity zone that hosts the DBMS and a privacy zone that hosts operators, with a secure channel preventing direct interface access by DBAs. For maintainability, HEDB records cross-zone calls in ciphertext, enabling DBAs to replay and debug queries without plaintext access. Fedb inherits and improves this design, but is designed to work with any modern CDB.
2.3 What Makes Modern CDB Systems Slow?
Modern CDBs exemplified by HEDB [15] have demonstrated significant performance overheads. Benchmarks on TPC-H show that HEDB can incur up to a 79.5 increase in latency compared to a plaintext database. These overheads render modern CDBs impractical for latency-sensitive applications, hindering broader industry adoption. Our analysis attributes them to two main factors.
Factor-#1: expensive en/decryption. In modern CDBs, each computation on sensitive data requires a cross-domain round-trip, involving two decryptions and one encryption, each taking thousands of CPU cycles (see Table 1). For an operation like SUM on a table with rows, the computation incurs round-trips, resulting in decryptions and encryptions. Given that N is often tens of millions in real-world scenarios (reported by several CDB vendors we surveyed), this high frequency of expensive en/decryptions inevitably introduces high overhead on the critical path. While batching multiple RPC invocations into a single round trip (e.g., for aggregation operators such as SUM/AVG/MAX/MIN) can help amortize overhead, the core issue remains that cryptographic operations must still be performed on every single data item.
Factor-#2: ciphertext expansion. Modern CDBs employ field-level encryption with AES-GCM. To guarantee data confidentiality and integrity, each encrypted field requires a nonce (12 bytes) and message authentication code (MAC, 16 bytes), causing ciphertext expansion: a 4-byte integer expands to 32 bytes, an 8 size increase. This expansion creates two performance bottlenecks. First, it enlarges the data transferred during frequent inter-domain RPCs, resulting in higher memory copy latency. Second, it increases the overall storage footprint of the database, leading to more I/O operations between the DBMS buffer pool and storage.
Empirical analysis. We profile HEDB [15] (commit f48742de) using three workloads: (i) TPC-H Q1 with a dataset larger than memory, representing large table computation; (ii) TPC-H Q1 with a dataset fitting in memory, capturing lightweight analytical settings; and (iii) a transactional benchmark with many write-heavy small queries. Figure 2 presents the breakdown of execution latency and storage costs. Despite the batching optimization, HEDB exhibits significantly higher overhead compared to the plaintext database. As shown in Figure 2 (a), the en/decryption operations, along with I/O amplification overhead, contribute considerably to the overall execution time, accounting for 10.5-62.6% and 13.9-25.5% across different workloads. Combined, they cause an overall latency slowdown of up to 5.3 compared to plaintext. Figure 2 (b) also shows that the field-level encryption incurs 1.4 to 3 storage overhead compared to plaintext. The remaining 8.6-26.5% of the execution time is attributable to DBMS execution and RPC overhead across domains, the latter of which stems from frequent copying of expanded ciphertexts.
Strawman: a ciphertext-to-plaintext cache. Since confidential VM-based TEEs (e.g., Intel TDX [19], AMD SEV-SNP [17]) provide sufficiently large secure memory, one can maintain a ciphertext-to-plaintext cache within TEEs to eliminate cryptographic operations from the critical path. However, this has two drawbacks. First, ciphertext search costs can be substantial: our preliminary experiment using C++’s STL unordered_map with one million ciphertexts showed an average lookup time of 1,601 CPU cycles, exceeding the constant-time decryption cost of 1,408 cycles. Second, storing both ciphertexts and plaintexts does not address the overhead of ciphertext expansion.
3 Fedb Overview
Design goals. Fedb has two design goals:
G1: Compatibility. Fedb aims to be compatible with the modern CDB design with functionality (standard SQL and ACID), maintainability (enabling DBA operations) and security (data confidentiality and database integrity).
G2: Efficiency. Fedb aims to reduce the performance gap between existing modern CDBs and the plaintext baseline, by optimizing two factors as pointed out in § 2.3.
Threat model. We assume a powerful adversary who controls the entire cloud infrastructure except TEE hardware. This includes compromised server software, rogue cloud and database administrators, and external attackers. Adversaries can read and modify any data outside TEE domains—data at rest, in untrusted memory, and in transit—and observe all network traffic and I/O patterns between TEE memory and persistent storage. We do not consider database vulnerabilities [33], TEE defects [34], denial-of-service or physical attacks.
We assume TEEs provide two essential properties: (i) memory isolation preventing external access to TEE-internal state, and (ii) remote attestation enabling users to verify code integrity before establishing trust. While TEE-agnostic, Fedb requires these guarantees to ensure that sensitive plaintext data remains visible only within TEE domains (§ 8).
Insights. Current practice of modern CDBs employs a mapping-based protection scheme, using ciphertexts as pointers to plaintext data in TEEs. This design conflates two orthogonal concerns: indirection (cross-domain references) and protection (data confidentiality). Because ciphertexts serve both roles, this coupling necessitates expensive en/decryption on every RPC. Fedb decouples them: FIDs provide efficient indirection, while encryption protects data at rest.
In fact, separating indirection from protection is a long-standing systems principle [35]. Capability systems [36, 37] used unforgeable tokens protected by OS confinement, not cryptography. Similarly, OSes have long used opaque handles, e.g., Unix file descriptors [38, 39], that are mapped internally by the kernel. In both cases, security comes from address-space isolation, keeping references simple and fast.
Architecture. Fedb’s high-level architecture is depicted in Figure 3. Fedb inherits HEDB’s dual-zone architecture but fundamentally redesigns how data flows between zones. The integrity zone runs a commodity DBMS engine, and the privacy zone comprises four components: (1) a proxy that performs mapping translations and invokes the appropriate operators, (2) expression operators that perform computation over plaintext secrets, (3) the mapping store that manages the FID-to-secret mappings, and (4) a write-ahead log (WAL) for ensuring cross-domain consistency. The mapping store is divided into temporary and permanent partitions for fine-grained management. To ensure data durability and transaction consistency, Fedb uses a commit protocol and a WAL for crash recovery. The proxy interacts with the mapping store through a set of well-defined APIs.
APIs. Fedb provides five APIs to manage mapping store:
Put(partition_num, secret) FID: Stores the given secret in the specified partition of the mapping store. Returns an FID with the partition number embedded, as a reference to the stored secret.
Get(FID) secret: Retrieves the secret associated with the specified FID, or returns a null value if the FID has no corresponding mapping.
Delete(FID): Removes the FID-to-secret mapping in the mapping store corresponding to the specified FID.
FlushLog(): Persistently flushes the mapping store’s WAL in a blocking I/O manner.
Recover(): Replays the WAL to restore the state of the mapping store after a crash.
With minor modifications using the provided APIs, existing modern CDBs [10, 12, 14, 13, 15] can enjoy the benefits of Fedb’s performance gains (see 1).
Workflow. Fedb’s query execution follows modern CDBs (described in § 2.2). To perform computation, the proxy fetches secrets associated with each input FID from the mapping store via Get(), invokes the appropriate expression operator to compute over these retrieved secrets, and then stores the result back to the mapping store via Put(), which generates a new FID representing the output. The DBMS proceeds with query execution using this returned FID. Finally, the proxy translates FIDs back into ciphertexts by calling Get() followed by encrypting the retrieved results within the privacy zone before returning them to users. This design preserves compatibility (G1) with modern CDBs that employ client-side encryption [12, 14, 13, 15].
Challenges. Since Fedb introduces states into the privacy zone, it brings two challenges. First, the lookup latency in the mapping store may grow with dataset size and eventually exceed constant-time cryptographic operations (§ 4). Second, maintaining state consistency between the integrity and privacy zones is necessary for database correctness (§ 5).
4 Achieving Efficiency via Mapping Store
To achieve efficiency (G2), Fedb addresses the two performance bottlenecks of modern CDBs by introducing the mapping store to implement crypto-free mappings.
4.1 Mapping Management
Data structure design. The mapping store is a linear array of tightly packed fields. Each sensitive field is assigned a unique integer as the FID that serves as a direct array index, hiding the size of individual fields from the adversary. This design restricts the metadata overhead per sensitive field to a mere 8B (FID), representing a 75% reduction compared to the 32B metadata required by modern CDBs. Compared to tree- or hash-based key-value stores (e.g., LevelDB [40], RocksDB [41], LMDB [42]), the design provides two advantages. First, it enables access complexity to sensitive fields in the privacy zone, eliminating the need for tree traversals or hash table probing. Second, synchronization is only necessary during FID allocation, which significantly reduces lock contention.
Mapping store operations. For fixed-length types (e.g., FLOAT8, TIMESTAMP), the Delete() operation marks the corresponding mapping entries as logically deleted. When inserting with Put(), Fedb either overwrites logically deleted slots or appends the new data to the end. For variable-length types (e.g., VARCHAR, NUMERIC), the situation is subtle because values cannot be overwritten in place. We address this by maintaining two arrays within the privacy zone: an offset array and a slab-like data array. The offset array is indexed directly by FIDs and stores pointers to the actual field. The data array is inspired by the slab allocator [43]: fields are grouped into size-class buckets, and each bucket is maintained as a contiguous packed array. This minimizes memory fragmentation for frequent allocation and deallocation. Operations within a single bucket follow the same overwrite/append logic as the fixed-length case.
Partition-based state management. The mapping store cannot directly infer the lifetime of each field. For example, when an expression operator produces a result, it is unclear whether this value will be discarded shortly after use or eventually persisted. Careless reclamation may leave live FIDs pointing to reclaimed storage, causing data loss.
To address this, we categorize values into two classes: (i) ephemeral values (e.g., query constants and intermediate results), and (ii) persistent values (e.g., data that survives query completion). These classes require distinct persistence and reclamation strategies, so we map them to separate mapping store partitions: a temporary partition for ephemeral values and a permanent partition for persistent values.
Ephemeral values are first placed in the temporary partition via Put() with the temporary partition identifier TMP_PARTITION_NUM. If the DBMS later decides to persist such a value, Fedb intercepts the write path and performs another Put() with the permanent partition identifier. During this Put() call, Fedb allocates a new persistent FID, copies the value into the permanent partition, and returns the new FID to the DBMS. At the end of each query, any values remaining in the temporary partition—now either intermediate results or temporary copies—can be safely discarded.
For persistent values, reclamation is triggered when the DBMS instructs the proxy to invoke Delete() on mappings associated with specific FIDs. This occurs when the DBMS needs to reclaim obsolete fields, for example, when PostgreSQL runs VACUUM to remove dead tuples.
4.2 Exploiting Data Locality
Mismatches in data access patterns between the DBMS and mapping store can cause major page faults on the critical path. For example, when the database accesses an FID whose corresponding secret is not in memory, Fedb must perform an extra I/O in the privacy zone, thereby increasing latency. To reduce such latency, we align the data layout between the DBMS and the mapping store, and proactively prefetch secrets likely to be accessed by the DBMS. This is achieved by exploiting both temporal and spatial locality.
Exploiting temporal locality. In Fedb, the DBMS must first retrieve the FID from storage, and then resolve it through the mapping store to obtain the secret. This dependency creates a natural alignment in temporal access patterns between the DBMS and the mapping store. For instance, when a query processes fields associated with FID1, FID2, and FID3 in order, the mapping store will access the same sequence to retrieve secret1, secret2, and secret3.
Exploiting spatial locality. To further improve performance, we divide permanent partitions to align the mapping store layout with the storage layout of the underlying database. The partitioning policy is customizable depending on the database in use. For example, in PostgreSQL, each table is a separate heap file with unordered field placement and reclamation. Fedb assigns one mapping store partition per heap file to preserve field colocation. In InnoDB, data resides in globally ordered fixed-size pages (e.g., 16 KB) with unordered fields inside each page. Fedb assigns one partition per page to maintain this alignment. We encode the partition number in the high bits of each FID to segment the FID space.
Prefetching. A random lookup in Fedb may involve two serial I/O operations: one for the DBMS to fetch FIDs and another for the mapping store to retrieve secrets. Fedb reduces this overhead by maximizing I/O overlap. Specifically, when the DBMS issues a disk read, Fedb intercepts it and identifies the partition being accessed. Since Fedb aligns accesses across the two domains at the granularity of partitions, it proactively sends an RPC to the proxy specifying which partition will be accessed. The mapping store then prefetches the corresponding partition concurrently with the DBMS’s I/O operation.
5 Ensuring Cross-domain Consistency
Modern CDBs rely on DBMSes to provide ACID properties. In contrast, Fedb executes transactions across two isolated domains, requiring new mechanisms that not only preserve ACID semantics but also guarantee cross-domain consistency.
We adopt external synchrony [44], a consistency model that distinguishes between internal system state (invisible to users) and external state (user-visible). External synchrony requires only that the external state remains consistent—internal states may temporarily diverge as long as they never violate external invariants. In Fedb, the mapping store constitutes internal state (users never see it), while the FIDs stored in the database represent external state. Therefore, external synchrony requires maintaining the invariant that every FID in the database maps to a valid, up-to-date secret in the mapping store. Temporary "orphan" secrets in the mapping store without corresponding database FIDs are permissible, as they remain invisible to users. To uphold this invariant, Fedb introduces lightweight synchronization between the DBMS and mapping store at transaction boundaries. We now detail the mechanisms for handling aborts, commits, and crash recovery.
Handling transaction aborts. When the DBMS encounters exceptions, all modifications made by aborted transactions must be correctly rolled back. To ensure rollback correctness, Fedb guarantees that all old versions of FIDs have corresponding secrets in the mapping store.
To this end, Fedb employs a multi-version concurrency control (MVCC)-like approach: old values from aborted transactions (e.g., deleted entries and values prior to updates) are logically deleted at runtime but physically reclaimed only when the DBMS garbage-collects them (e.g., PostgreSQL’s off-path VACUUM). For instance, as illustrated in Figure 4 (a), when a transaction issues an update, Fedb first uses Delete() to mark the original value as logically deleted (①), then inserts the new value out-of-place (②). If the transaction is aborted, the DBMS can access the original values as usual (③). The old value is eventually reclaimed by the DBMS’s offline garbage collection process (④).
Handling transaction commits. While the above mechanism maintains the invariant at runtime, a privacy-zone crash can cause the mapping store to lose recently written secrets even though their corresponding FIDs are already persisted by the DBMS, violating the requirement of external synchrony.
Fedb fixes this by enforcing a strict persistence order: the mapping store must durably persist before the DBMS commits. As shown in Figure 4 (b), the commit protocol has three phases: prepare, commit #1, and commit #2. During the prepare phase, Fedb intercepts the DBMS’s WAL flush events. In the commit #1 phase, Fedb invokes FlushLog() to persist its own WAL. Only after this step completes does the DBMS persist its log in the commit #2 phase, ensuring that secrets are durable before their corresponding FIDs become visible to users.
Recovering from system crashes. We first discuss privacy-zone crashes. Upon detecting an RPC timeout, the DBMS aborts the affected transactions and sends a reboot signal to the privacy zone. After the privacy zone reboots, the proxy starts a WAL replay. Once the privacy zone has successfully recovered, Fedb resumes normal operation.
Next, we consider integrity-zone crashes, where the DBMS performs its own recovery using standard techniques. Our commit protocol ensures the external synchrony invariant remains satisfied, but this may sometimes lead to storage overhead. For example, if the integrity zone crashes while waiting for commit confirmation from the privacy zone (the ② crash point in Figure 4 (b)), some secrets remain recoverable from the privacy zone’s WAL, but their corresponding FIDs in the DBMS are lost. To identify and reclaim these “orphaned” secrets, Fedb periodically performs garbage collection.
6 Implementation
We implemented Fedb in 5.8K lines of C++ code on PostgreSQL 15.5. Table 2 lists the code complexity of each component. Fedb leverages ARMv8.4 Secure EL2 [45, 24] to create two isolated confidential VMs as integrity and privacy zones. Both use dm-integrity [46] for corruption detection; the privacy zone additionally applies dm-crypt [47] for transparent I/O encryption. The mapping store is implemented as memory-mapped (mmap-based) arrays for persistence. Temporal locality is handled by an LRU cache combined with Linux sequential prefetching, while spatial locality is achieved by aligning partitions with database tables. Each 64-bit FID encodes a configurable 16-bit partition prefix and a 48-bit offset; unique FIDs within a partition are generated via an atomic monotonic counter. Users can tune FID length and partition-prefix size according to expected data volume—shorter FIDs reduce storage overhead and improve lookup efficiency.
| Fedb Components | Language | LoC |
| mapping store | C++ | |
| WAL | C++ | |
| Commit Protocol (in PostgreSQL source) | C | |
| Proxy | C++ | |
| Operators (from HEDB [15]) | C++ | |
| Crypto Library | C |
7 Performance Evaluation
We evaluate Fedb by answering the following questions:
Experimental setup. We evaluate Fedb on a Secure-EL2 enabled ARM server machine with a 128-core CPU (2.9 GHz, 4 NUMA nodes), 512 GB DDR4 DRAM and 1 TB SSD running Ubuntu 22.04 LTS. The experiments are performed on 2 confidential VMs running Linux 6.2.0. The integrity zone runs with 32 vCPUs and 64 GB of memory, whereas the privacy zone runs with 16 vCPUs and 32 GB of memory. Both disable memory swap. To reduce cross-NUMA RPC cost and exclude hypervisor scheduling overhead, we fix their CPU affinity in the same NUMA node.
7.1 End-to-end Benchmarks (RQ1)
We use two standard benchmarks: TPC-C and TPC-H. For comparison, we include the plaintext database (original PostgreSQL) and a state-of-the-art, HEDB [15].
TPC-C. Following Microsoft Always Encrypted [12], we encrypt personally identifiable columns (e.g., C_STREET_1, C_CITY, C_STATE in the Customer table). We evaluate TPC-C throughput (transactions per second, or TPS) across varying numbers of connection threads using 128 warehouses. To prevent data-size leakage, we pad variable-length types (e.g., VARCHAR) to a fixed length of 128 B. After a 120-second warm-up, we measure for 600 seconds. As shown in Figure 5, throughput peaks at 8 connections and then degrades due to contention. Across all bars, Fedb consistently outperforms HEDB. HEDB achieves only 51.6–77.3% of baseline throughput, whereas Fedb reaches 63.3–91.3%, delivering up to 1.8 speedup.
TPC-H. Following Alibaba HEDB [15], we use TPC-H (scale factor = 3), a standard OLAP benchmark with large joins, group-bys, aggregations, and arithmetic operations. We make all non-key columns sensitive, as TPC-H does not define security attributes. Both Fedb and HEDB can enable batching (with a batch size of 256 fields), which aggregates multiple operations into a single RPC instead of iteratively. We also report results with batching disabled. Results are averaged over 5 runs per query after one warm-up run.
Figure 6 shows the slowdown of HEDB and Fedb relative to plaintext PostgreSQL. With batching enabled, HEDB incurs 22.7 average slowdown, driven by excessive cryptographic operations and I/O amplification due to ciphertext expansion. In contrast, Fedb limits the slowdown to 0.9-13.3 (3.7 on average) and delivers up to 78.0 speedup over HEDB. Disabling batching degrades performance for both systems, since aggregation operators like SUM/AVG/MIN/MAX can no longer process data in bulk. Even without batching, Fedb remains consistently faster than HEDB. Interestingly, Fedb slightly outperforms the plaintext PostgreSQL on Q3. This occurs because Fedb stores only fixed-size FIDs, enabling PostgreSQL’s query planner to apply more aggressive and efficient physical optimizations than when handling variable-length data types in plaintext.
(a) Transaction Modes (Zipfian, factor=0.8)
(b) Transaction Modes (Uniform)
(c) Total Memory Size
Storage consumption. As shown in Table 3, Fedb reduces storage consumption by 34.3–80.0% compared to HEDB across evaluated workloads. This is because Fedb adds an 8-byte FID per value, while existing modern CDBs attach 28 bytes of cryptographic metadata per field. In addition, HEDB requires padding each sensitive field to an upper bound to hide sizes (if necessary), which further amplifies I/O and memory pressure. In contrast, Fedb by design hides the size information as it replaces all fields with constant-size FIDs.
| Database | TPC-H (S=3) | TPC-C (W=128) |
| PostgreSQL | 5.2 | 13.3 |
| Fedb | 8.4 (1.6) | 15.1 (1.1) |
| HEDBw/o | 17.8 (3.4) | 20.4 (1.5) |
| HEDBw | 42.1 (8.1) | 23.0 (1.7) |
w/o: without plaintext padding w: with plaintext padding
7.2 Sensitivity Analysis (RQ2)
The performance gain of Fedb depends on both workload characteristics and system configuration. To characterize the impact of these factors, we use Sysbench [48] to issue transactional microbenchmarks with configurable read/write ratios and access patterns. The benchmark runs on a synthetic dataset consisting of 32 tables with 1M rows each, occupying approximately 7 GB, 10 GB, and 12 GB in plaintext PostgreSQL, Fedb, and HEDB, respectively. To control for other variables, we set the total available memory equal to the machine’s physical memory, except in the memory-sensitivity experiments. When experimenting with memory sensitivity, we use cgroups to adjust the total available memory (the sum of the integrity and privacy zones) to match different sizes of the dataset. Each experiment runs a 120-second warm-up followed by a 600-second measurement, using 16 client connections by default.
Transaction sensitivity. Figure 7 (a) presents normalized throughput under Zipfian data distribution (factor = 0.8) across six operation modes: read-only, read-write, write-only, insert-only, point-select, and range-select. Fedb achieves 1.1-5.8 speedup over HEDB and delivers 46.6-95.0% of plaintext throughput. In write-only mode, Fedb is 4.6 faster than HEDB. This gap arises from frequent UPDATE queries such as “UPDATE test SET k=k+1;” where Fedb avoids the costly decrypt–compute–re-encrypt cycle required by HEDB.
Skew sensitivity. We evaluate against a random access pattern generated from a uniform distribution, which exhibits relatively poor locality. As illustrated in Figure 7(b), even under a non-skewed access pattern that yields frequent CPU cache and TLB misses, Fedb still outperforms HEDB by up to 2.3 across all six modes, thanks to its crypto-free mapping.
Memory sensitivity. Both HEDB (with ciphertext expansion) and Fedb (with FIDs overhead) demand more memory than plaintext PostgreSQL, making them more sensitive to available memory size than their plaintext counterpart. Memory constraints impact HEDB and Fedb differently. In contrast, HEDB suffers from much more severe storage expansion (as shown in Table 3). Fedb’s overall performance is directly related to the average mapping lookup latency; limited memory will trigger frequent major page faults on the critical path.
To quantify this sensitivity, we measure their performance under different total memory sizes. We use a read-write workload with 8 threads and a random access pattern. As illustrated in Figure 7(c), Fedb consistently outperforms HEDB, with an average speedup of 1.9 and a 2.2 speedup even when the total available memory is limited to 1 GB (10% of the dataset for Fedb). This is because Fedb has a smaller storage footprint than HEDB and preserves the same locality between the mapping store and the database.
Technique sensitivity. Under the read-write mode with 16 threads, we evaluate the impact of three key techniques: (1) temporary partitioning (§ 4.1), (2) permanent partitioning (§ 4.2), and (3) transaction coordination (§ 5). Temporary partitioning improves throughput by 2.4%, with faster FID allocation via free lists outweighing extra RPC cost during persistence. Permanent partitioning delivers the highest gain (+7.0%) by maximizing data locality. Transaction coordination incurs negligible overhead (-0.90%), as each transaction commit triggers a lightweight FlushLog() call.
7.3 Miscellaneous (RQ3)
System recovery cost. We run Fedb for 5 minutes in microbenchmark read-write mode, then abruptly terminate its process to simulate a crash (generating 303 MB of WAL) and restart. The WAL replay in the mapping store takes 2.1 seconds, whereas the DBMS needs 13.2 seconds. When both zones crash, their replays occur in parallel; the faster mapping store recovery is fully masked by the slower DBMS recovery. If only the privacy zone crashes, the DBMS in the integrity zone pauses for 2.1 seconds. To shorten this pause, Fedb supports a configurable WAL size bound, which allows for trading runtime performance for faster recovery: when reaching this bound, Fedb triggers earlier log truncation.
Performance on Intel x86. Our evaluation is based on ARM commodity hardware, as it relies on ARM TEE features [24]. We also measure operation costs on an Intel Xeon x86 platform that supports AES-NI hardware acceleration. The results on x86 show that en/decryption remains an order of magnitude slower than put/get operations, as detailed in Table 4.
| Encrypt | Decrypt | Put | Get | |
| ARM | 5795 | 1408 | 55 | 30 |
| x86 | 1256 | 509 | 76 | 50 |
8 Security Analysis
Consistent with prior modern CDBs [10, 12, 14, 13, 15], we consider two adversary classes: active adversaries can arbitrarily tamper with storage and modify untrusted memory, and passive adversaries (honest-but-curious) do not modify data but attempt to infer sensitive information through observation.
Defense against active attacks. Fedb ensures that sensitive data remains confidential and fresh. First, TEE hardware provides memory encryption and isolation, preventing adversaries from accessing plaintext data during computation. Second, when data is evicted to untrusted storage, Fedb encrypts it using authenticated encryption at the block level with a counter-based freshness mechanism [49], which can detect storage corruption and rollback attacks [50, 51].
Fedb ensures faithful query execution through two mechanisms. First, remote attestation allows users to verify legitimate Fedb code is running in the TEEs before uploading data. Second, following HEDB [15], Fedb places both the DBMS (in the integrity zone) and expression operators (in the privacy zone) within trusted domains, with secure channels between TEEs that prevent message injection, replay, man-in-the-middle and interface attacks.
Defense against passive attacks. The primary question is whether an adversary observing FIDs can learn anything about the corresponding plaintext secrets. We prove that Fedb provides indistinguishability. The observation is that FID allocation is stateful but plaintext-independent. The allocator maintains an internal counter (tracking the next available FID) and assigns FIDs deterministically from the counter, independent of the plaintext . After prior Put() calls, any two plaintexts and submitted as the -th item receive the exact same FID. Thus, the adversary’s view of the FID sequence is identical for and . Fedb’s crypto-free mappings provide confidentiality comparable to semantic security [52]: no information about plaintexts is revealed (beyond leakage from other channels, discussed later).
Additionally, Fedb provides collision resistance: different plaintexts receive different FIDs due to the large FID space ( possible values). Even inserting the same plaintext yields distinct FIDs, mirroring the behavior of non-deterministic encryption schemes like AES-GCM.
Leakage analysis. While Fedb protects plaintext contents, practical query functionality requires revealing certain metadata to the untrusted infrastructure. We characterize Fedb’s leakage profile using the definition of -security [53, 10], which models security as a leakage function that precisely specifies what information the adversary learns. We define Fedb’s as follows:
: During database initialization with dataset , the adversary learns the total number of fields from untrusted storage. However, since all fields are replaced by constant-size FIDs, individual record sizes remain hidden.
: For query over , the adversary learns:
1.
I/O access pattern : which storage blocks (indexed by FIDs) are accessed during query execution.
2.
Search pattern : whether two queries access overlapping record sets, revealing query repetition.
3.
Timing : when and how long the query executes.
4.
Result size : the number of records returned or affected by the query.
- 5.
1.
: For update operation (i.e., INSERT, UPDATE, DELETE), the adversary learns which storage blocks are modified and the operation type, but not the plaintext values being written or deleted.
This leakage profile provides -semantic security: an adversary observing Fedb’s execution learns nothing beyond about the underlying plaintext data.
Side-channel considerations. Fedb’s stateful privacy zone retains plaintext data in TEE memory for extended periods (until eviction to storage). This residency could increase the attack window for digital side-channels (e.g., cache timing or memory access patterns). Modern CDBs already face similar risks, as encryption keys reside permanently in TEE memory—key compromise would breach all sensitive data. Existing mitigation for key protection (e.g., strict CPU resource isolation [54, 55]) can equally safeguard plaintext data in TEEs. To protect against controlled-channel attacks [56], Fedb utilizes ARM Secure EL2 [45] that fully isolates TEEs and their page tables from untrusted privileged software. Importantly, Fedb’s design is generalizable to other TEEs offering equivalent guarantees, including Intel TDX [19], AMD SEV-SNP [17], IBM PEF [25], etc. Physical attacks that infer DRAM access patterns remain prohibitively expensive [57].
9 Related Work
Confidential databases (CDBs). To address rising privacy concerns, CDBs on untrusted clouds have long been a goal [58]. Systems like TrustedDB [7], EnclaveDB [9], DBStore [32], and SecuDB [16] shield an entire DBMS in TEEs for strong security with little maintainability. Modern CDBs, like AlwaysEncrypted [12], GaussDB [14], Operon [13], and HEDB [15] prioritize maintainability, yet often incur severe performance overhead. Motivated by the growing industrial adoption of this latter class of systems, Fedb aims to improve their efficiency while preserving their advantages.
Other CDB systems take different approaches: fully cryptographic databases [59, 53, 60, 61, 62, 63, 64, 65] require no trusted hardware, and oblivious databases [66, 67, 68] hide access patterns. These lines of work are orthogonal to ours.
Indirection-protection separation. Separating indirection from protection is a classic systems principle [35]. Capability systems [36, 37] pioneered the use of unforgeable tokens (capabilities) as references to resources protected by OS isolation. Similarly, OSes have long used opaque handles [38, 39]: file descriptors in Unix are integers mapped by the kernel to internal objects. Fedb shares this insight: data-independent FIDs serve as capabilities protected by TEE isolation, removing cryptographic overhead on every reference.
Applicability beyond CDBs. Distributed confidential analytics systems, like Spark-based Opaque [69] and Laputa [70], Hadoop-based Civet [71] and VC3 [72], encrypt data during cross-node shuffles, incurring costly cryptographic operations at TEE boundaries. These systems may benefit from the abstraction of crypto-free mappings for better efficiency.
10 Conclusion
This paper challenges a core assumption in modern CDB design: mappings between untrusted and trusted domains must be cryptographic. Existing CDB systems using ciphertext-to-plaintext mappings suffer from critical performance bottlenecks. Fedb introduces crypto-free mappings, which eliminate costly en/decryption from the critical path and avoid ciphertext expansion. Our deployment in production DBaaS systems validates the practicality of this approach.
Acknowledgments
This work was sponsored by CCF-Huawei Populus Grove Fund. Mingyu Li is the corresponding author.